Circular No. 8714

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 K2 (McNAUGHT) R. H. McNaught reports his discovery of another comet with a 20′′

coma on Uppsala Schmidt images taken in poor seeing (discovery observation tabulated below). Better images obtained by McNaught in good seeing on May 23.6 UT show a strong central condensation and a weak coma with a diameter of 20′′, extended slightly in p.a. 300^{o}. Following posting on the ‘NEO Confirmation Page’, R. Trentman and R. Fredrick (Louisburg, KS, 0.75-m reflector) report that their CCD images obtained on May 23.2–23.3 show a diffuse coma several arcsec in diameter.

2006 UT  α_{2000}  δ_{2000}  Mag.
May 22.64208  15^{h}33^{m}16.^{s}38  −28^{o}22′37′′.4  17.7

The available astrometry, the following preliminary orbital elements, and an ephemeris appear on MPEC 2006-K32.

T = 2006 Aug. 3.842 TT  ω = 256.^{o}183
e = 0.49732  Ω = 13.158 }2000.0
q = 2.02900 AU  i = 6.468

a = 4.03634 AU n^{o} = 0.121541 P = 8.11 years

COMETS C/2006 F5–F8 AND C/2006 G2 (SOHO) Additional comets have been found on SOHO website images (cf. IAUC

8713; MA = M. Mazzucato), all Kreutz sungrazers except C/2006 F6 (Meyer group). C/2005 F5 was described by K. Battams as small, slightly diffuse, and too faint for photometry. C/2006 F6 was condensed and quite elongated, reaching mag 7.6 on Mar. 22.993 UT at 7.1R_{⊙}. C/2006 F7 and C/2006 G2 were extremely faint and diffuse. C/2006 F8 reached mag 4.7 in C3 images on Mar. 30.571 at 12.0R_{⊙}; in C2 images it showed a very faint, thin tail that reached a length of 21′ on Mar. 30.922 at 5.1R_{⊙}, with a possible secondary component (‘lump’ on the side of the head).

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 F5 Mar. 21.321 0^{h}18.^{m}4 − 0^{o}33′ C3  MA 2006-K10
C/2006 F6  22.963 0 12.2 + 2 37 C2  TH 2006-K10
C/2006 F7  28.326 0 36.1 + 2 34 C2  RM 2006-K10
C/2006 F8  29.654 1 00.3 + 1 58 C3/2 HS 2006-K10
C/2006 G2 Apr. 11.246 1 27.0 + 7 33 C2  TH 2006-K29
2006 May 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green